# Continuous-time multifarious systems - Part II: non-reciprocal multifarious self-organization

### v2

@@ D.1 Nucleation calculation
$\frac{R_{\mathrm{sq}}}{R_{\mathrm{sq}^{\prime}}}=e^{3\lambda-2\varepsilon}.$ (22)
@@
For $\lambda>\varepsilon$ the transition that dominates is $\mathcal{A}_{N}$, which has rate $R_{\mathcal{A}_{N}}=e^{\mu+4\lambda-4\varepsilon/2}$. The dynamics evolve via many individual nucleations of isolated $S^{2}$ tiles. The time to reach some small fraction $f$ of $S^{2}$ in the system is thus given by
$t_{N}({\lambda>\varepsilon})$ $\approx fL^{2}\frac{1}{L^{2}R_{\mathcal{A}_{N}}+\cdots},$ (23)
$\approx\frac{f}{R_{\mathcal{A}_{N}}},$ (24)
@@
The dots in the first line represent the contribution to the total rate at each step from other transitions. In this parameter range the other transitions have a negligible rate compared to the $\mathcal{A}_{N}$ transition.
@@
$t_{N}\left({\frac{2}{3}\varepsilon<\lambda<\varepsilon}\right)$ $=\frac{R_{\mathcal{A}^{\prime}_{N}}}{R_{\mathcal{B}_{N}}}\left(\frac{1}{L^{2}R_{\mathcal{A}_{N}}}+\frac{1}{R_{\bar{\mathcal{A}}_{N}}}\right)+\cdots,$ (26)
$\approx\frac{R_{\mathcal{A}^{\prime}_{N}}}{R_{\mathcal{B}_{N}}}\left(\frac{1}{L^{2}R_{\mathcal{A}_{N}}}+\frac{1}{R_{\mathcal{A}^{\prime}_{N}}}\right),$ (27)
$=\frac{1}{L^{2}}e^{-\mu-7\lambda+4\varepsilon}+e^{-\mu-4\lambda+2\varepsilon}.$ (28)
Here the dots in the first line represent the time to grow after a seed has successfully reached the critical point, which is negligible compared to the time required to get a seed to the critical point. Here we see a crossover in scaling from
$t_{N}\left({\frac{2}{3}\varepsilon\lesssim\lambda<\varepsilon}\right)\sim e^{-7\lambda},$ (29)
for $\lambda$ close to but greater than $\frac{2}{3}\varepsilon$, to
$t_{N}\left({\frac{2}{3}\varepsilon<\lambda\lesssim\varepsilon}\right)\sim e^{-4\lambda},$ (30)
for $\lambda$ close to and below $\varepsilon$, and any value above $\varepsilon$ as derived in Eq. (25).
@@ D.2 Interface growth calculation
To calculate the interface velocity, we start by calculating the average time taken to grow a full layer. Consider an initially flat interface. The first step is to nucleate the next layer, which is done in the transition labeled $\mathcal{A}_{I}$. The rate of the reaction to do this is $R_{\mathcal{A}_{I}}=e^{\mu+3\lambda-2\varepsilon/2}$. For $\lambda<\varepsilon$ this transition dominates over the $\mathcal{B}_{I}$ transition, and so the average time for this step is $t_{\mathcal{A}_{I}}\approx(LR_{\mathcal{A}_{I}})^{-1}$. The factor of $L$ is present since there are $L$ different sites along the interface where this reaction could occur. Now the dominant transitions are to grow out this new layer, using transitions like $\mathcal{C}_{I}$. This involves $L-2$ reactions, each of which takes on average a time $t_{\mathcal{D}_{I}}\approx(2e^{\mu+2\lambda})^{-1}$. Adding the final $S^{2}$ tile to the new layer has a different rate, since the site is now surrounded by two $S^{2}$ tiles. This step takes a time $t_{\mathcal{E}_{I}}\approx(e^{\mu+2\lambda+2\varepsilon/2})^{-1}$. Putting this all together, the total average time to grow a new layer is
@@
To identify the upper bound in $\lambda$, one can see that for $\lambda>\varepsilon$, direct replacement of $S^{2}$ tiles in the bulk of $S^{1}$ becomes the dominant reaction, as discussed in the calculation of the nucleation time. This leads to the breakdown of interface growth, and so we consider the interface velocity to be zero in this parameter range.

### v3

@@ D.1 Nucleation calculation
$\frac{{\mathcal{R}}_{\mathcal{B}_{N}}}{{\mathcal{R}}_{\mathcal{B}^{\prime}_{N}}}=e^{3\lambda-2\varepsilon}.$ (22)
@@
For $\lambda>\varepsilon$ the transition that dominates is $\mathcal{A}_{N}$, which has rate ${\mathcal{R}}_{\mathcal{A}_{N}}=e^{\mu+4\lambda-4\varepsilon/2}$. The dynamics evolve via many individual nucleations of isolated $S^{2}$ tiles. The time to reach some small fraction $f$ of $S^{2}$ in the system is thus given by
$t_{N}({\lambda>\varepsilon})$ $\approx fL^{2}\frac{1}{L^{2}{\mathcal{R}}_{\mathcal{A}_{N}}+\cdots},$ (23)
$\approx\frac{f}{{\mathcal{R}}_{\mathcal{A}_{N}}},$ (24)
@@
The dots in the first line represent the contribution to the total rate at each step from other transitions. In this parameter range the other transitions have a negligible rate compared to the $\mathcal{A}_{N}$ transition. In discrete-time Monte Carlo simulations this parameter range is always unstable [5], destroying any structure in the system. In continuous-time Gillespie simulations, a shifting sequence $S^{1}\to S^{2}$, can successfully shape-shift. This is because the rates of the $\mathcal{A}_{N}$ transitions, and subsequent filling-in of the checkerboard pattern, are fast enough to shape-shift before the structure becomes unstable. With any other shifting sequence the Gillespie simulations do not lead to successful shape-shifting, because subsequent shifts start too quickly.
@@
$t_{N}\left({\frac{2}{3}\varepsilon<\lambda<\varepsilon}\right)$ $=\frac{{\mathcal{R}}_{\mathcal{A}^{\prime}_{N}}}{{\mathcal{R}}_{\mathcal{B}_{N}}}\left(\frac{1}{L^{2}{\mathcal{R}}_{\mathcal{A}_{N}}}+\frac{1}{{\mathcal{R}}_{\mathcal{A}^{\prime}_{N}}}\right)+\cdots$ (26)
$=\frac{1}{L^{2}}e^{-\mu-7\lambda+5\varepsilon}+e^{-\mu-3\lambda+\varepsilon}.$ (27)
Here the dots in the first line represent the time to grow after a seed has successfully reached the critical point, which is negligible compared to the time required to get a seed to the critical point.
@@ D.2 Interface growth calculation
To calculate the interface velocity, we start by calculating the average time taken to grow a full layer. Consider an initially flat interface. The first step is to nucleate the next layer, which is done in the transition labeled $\mathcal{A}_{I}$. The rate of the reaction to do this is ${\mathcal{R}}_{\mathcal{A}_{I}}=e^{\mu+3\lambda-2\varepsilon/2}$. For $\lambda<\varepsilon$ this transition dominates over the $\mathcal{B}_{I}$ transition, and so the average time for this step is $t_{\mathcal{A}_{I}}\approx(L{\mathcal{R}}_{\mathcal{A}_{I}})^{-1}$. The factor of $L$ is present since there are $L$ different sites along the interface where this reaction could occur. Now the dominant transitions are to grow out this new layer, using transitions like $\mathcal{C}_{I}$. This involves L2L-2 reactions, each of which takes on average a time $t_{\mathcal{C}_{I}}\approx(2e^{\mu+2\lambda})^{-1}$. Adding the final $S^{2}$ tile to the new layer has a different rate, since the site is now surrounded by two $S^{2}$ tiles. This step takes a time $t_{\mathcal{D}_{I}}\approx(e^{\mu+2\lambda+2\varepsilon/2})^{-1}$. Putting this all together, the total average time to grow a new layer is
@@
In practice a second layer can nucleate before the first one has finished growing - particularly for large interfaces and for $\lambda$ values closer to $\varepsilon$. These additional reactions are not included in the calculation, in the spirit of only considering the most dominant reactions.
To identify the upper bound in $\lambda$, one can see that for $\lambda>\varepsilon$, direct replacement of $S^{2}$ tiles in the bulk of $S^{1}$ (transition $\mathcal{B}_{I}$) becomes the dominant reaction, as discussed in the calculation of the nucleation time. This leads to the breakdown of interface growth, and so we consider the interface velocity to be zero in this parameter range.